A force 4N acts along the X- direction. Another force 5N makes an angle

60° with the first force. Find the magnitude and direction of the resultant.


1
Expert's answer
2021-02-18T18:39:40-0500
"R_x=4\ N\cdot cos0^{\circ}+5\ N\cdot cos60^{\circ}=6.5\ N,""R_y=4\ N\cdot sin0^{\circ}+5\ N\cdot sin60^{\circ}=4.33\ N."

The resultant of the sum of two forces can be found from the Pythagorean theorem:


"R=\sqrt{R_x^2+R_y^2}=\sqrt{(6.5\ N)^2+(4.33\ N)^2}=7.81\ N."

We can find the angle as follows:


"\theta=cos^{-1}(\dfrac{R_x}{R})=cos^{-1}(\dfrac{6.5\ N}{7.81\ N})=33.6^{\circ}."

The resultant R has magnitude 7.81 and angle of "33.6^{\circ}" above the positive x-axis.